# declare library=forked edges

### v1

@@ 1 Introduction
Large language models (LLMs), such as ChatGPT, Gemini, Claude, and Cohere [1, 2, 3, 4], have demonstrated remarkable capabilities in text generation, machine translation, and knowledge understanding tasks [5, 6, 7, 8, 9]. They effectively mimic human writing behaviors and generate complex and coherent outputs from the input text, making it challenging to determine whether a text is authored by humans or generated by LLMs. Due to the high demands of computational resources and human efforts required for training LLMs [10, 11], these models are commonly offered as a service through application programming interfaces (APIs), typically requiring users to pay or subscribe [12, 13]. Although users cannot access to the model weights or architectures of these commercial LLMs, this restriction does not ensure the safety of these models. Malicious actors can intentionally mimic cloud-hosted LLM behaviors to offer cheaper services [14, 15]. To conduct this service stealing, an adversary can query a set of inputs through an LLM’s API to retrieve the corresponding outputs. Then, the adversary uses these input-output data to effectively fine-tune their local model. When the number of queries is sufficient to gather enough input-output data within a particular domain, the adversary can steal the cloud-hosted LLM behaviors in that domain [16, 17, 18]. Consequently, these concerns underscore significant risks regarding the intellectual property (IP) rights of the cloud-hosted proprietary LLMs [19].
@@
To address such risks, service providers, i.e., the cloud, have employed many strategies including watermarks, encryption, limited model exposure via APIs, and differential privacy [9, 20, 21]. Among them, watermarks (WMs) [9, 22, 23, 24, 25] have emerged as a practical tool for LLMs due to its ability to ensure traceability, protect IP, and detect misuse. Typically, WMs embed imperceptible patterns into the LLM outputs directly. These patterns can be used to trace the original text or determine whether the generated text from an LLM is watermarked, facilitating the detection of unauthorized use. To watermark outputs, the cloud can introduce bias into logits of token generations to favor a set of specific tokens or alter the token sampling process of LLMs. Once an WM is applied, the cloud can use IP checkers to determine if a set of outputs from a suspicious model is watermarked. This allows the cloud to assess whether the suspicious model has been trained or fine-tuned using the cloud’s watermarked outputs. Typically, IP checkers analyze the tokens and perform statistical tests to check if the values exceed certain thresholds, indicating that the cloud-hosted LLM has been imitated.
@@
WMs are effective in detecting IP violations. However, WMs can be susceptible to WM removal attacks [26, 27] or spoofing attacks [27]. These attacks either require significant computational resources for additional training to paraphrase the generated text of cloud-hosted LLMs or demand numerous queries to learn the distribution of WMs. While several studies have surveyed WMs on LLMs [28, 29, 30], no research has systematically and extensively explored the effectiveness of WMs on LLMs, particularly regarding the impact of WMs on LLM utility and IP checkers, under both attack and attack-free environments. In addition, the practicality of WMs has not been rigorously assessed. Therefore, in this work, we aim to develop a comprehensive systematization of WMs in LLMs, with the following contributions:
@@ 2.1 Large Language Models
Large language models (LLMs) [31, 2, 32, 33] are advanced deep learning models, pre-trained on vast datasets with a large number of parameters, enabling them to perform a wide range of natural language processing (NLP) tasks, such as text classification, text and code generation, and question answering.
@@
In generation tasks, the model predicts each token in an autoregressive manner [31], similar to Eq. 1, to generate a complete text sequence.
The most widely used LLM architectures are based on transformers [34] and incorporate a self-attention mechanism. This mechanism includes a stack of multi-head attention and feed-forward layers. Positional encoding with distinct matrices for different positions ensures the model retains sequence information without relying on recurrent networks. Based on transformer architectures, existing LLMs can be divided into three main categories (Fig. 1), as follows. (1) Encoder-only models, such as BERT [35], RoBERTa [36], XLNet [37], or XLM [38], capture word context and are well-suited for tasks that require a deep understanding of input data, such as sentence classification, sentiment analysis, and information extraction. These models are trained bidirectionally, predicting masked words using both preceding and following words. (2) Decoder-only models, such as GPT and its family [39], operate unidirectionally, predicting the next token based on context in an auto-regressive manner. They are designed for text generation tasks without a separate encoding phase. Lastly, (3) Encoder-decoder models such as T5 [40], BART [41] or MASS [42] combine both encoder and decoder components for tasks that involve understanding and generating data, such as text summarization and translation.
@@ 2.2 Risks in LLMs
Due to the high demands of computational resources and human efforts required for training LLMs [10, 11], LLMs are commonly offered as a service through APIs, typically requiring users to pay or subscribe [12, 13]. Table VI (Appendix) presents a comprehensive overview of popular LLM APIs, highlighting their features, advantages, and constraints. According to the Artificial Analysis leaderboard [43], OpenAI and Microsoft are currently the leading providers in the LLM API market.
Although users may not have access to the model weights or architectures via APIs, this restriction does not guarantee the model’s safety. In addition, as LLMs become more widespread, the risk of their misuse for malicious purposes is increasing. Potential misuses of LLMs include several significant threats, as follows. First, LLMs can be exploited to generate false information or carry out harmful actions, such as crafting phishing emails, hate speech, or biased content to manipulate public opinion or amplify harmful narratives [44]. Second, LLMs can generate deepfake data, making manipulated media more convincing and enabling sophisticated scams or damage to reputation [45]. Third, adversaries can use LLMs to mimic individuals or organizations in text-based communications, facilitating theft, fraud, or financial crimes [46]. Fourth, LLMs can assist in writing malicious code or crafting evasion techniques to bypass detection systems, posing a significant cybersecurity threat [47]. More importantly, malicious actors may intentionally exploit LLM behaviors by imitating cloud-hosted models to offer cheaper services or engage in model-stealing attacks, further amplifying the aforementioned risks. Such attacks not only threaten the security of proprietary models but also expose users to fraudulent or unethical services [14, 15].
Figure 2: Model Stealing Attacks [48].
@@ 2.3 Model Stealing Attacks
In model stealing attacks, adversaries aim to imitate the behavior of a cloud-hosted model denoted as $\theta$, by constructing their local model $\theta_{adv}$ through fine-tuning on input-output pairs obtained by querying the cloud-hosted model [49]. The goal is to bypass IP checkers while maintaining the adversary model’s utility, which is comparable to that of the original model. To launch the attacks (Fig. 2), adversaries start with a set of $N$ task-specific prompts $\{x_{i}\}_{i=1}^{N}$, sending them to the cloud to obtain outputs $\{y_{i}\}_{i=1}^{N}$. These pairs are used to fine-tune $\theta_{adv}$ to mimic the behaviors of $\theta$.
@@
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, rectangle, rounded corners, align=left, text centered, minimum width=4em, edge+=gray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, , where level=1text width=10em,font=,, where level=2text width=10em,font=,, where level=3text width=11em,font=,, where level=4text width=7em,font=,, [Watermarking
for LLMs, leaf-head [WM During
Logits Generation, blue-box-lv2 [Mechanism: KGW [9], SIR [50], DiPmark [51], SemaMark [52], Adaptive WM [53],
Unbiased WM [54], TSW [55], WatME [56], GumbelSoft [57], MPAC [58], UPV [23], Unigram [59],
Topic-based WM [60], REMARK LLM [61], SWEET [62], CTWL [63], EWD [64], X-SIR [65],
SW [66], CodeIp [67], OW [68], Stylometric WM [69], NS-WM [70].
Strengths: Flexibility and non-intrusive WM, Effectively track watermarked text.
Weaknesses: Possible impact on semantic meaning, Vulnerable to removal attacks. , blue-box-lv4, text width=42.5em] ][WM During
Sampling Process, green-box-lv2 [Token-based
Sampling, green-box-lv3 [Mechanism: Undetectable WM [24], EXP [25], EDRW [71],
EXPGumbel [72], PDW [73], Bileve [74], STA-1 [75], SynthID [76].
Strengths: Can be incorporated easily and less noticeably, Simple
detection, checking the alignment between tokens and the sampling pattern.
Weaknesses: Can be vulnerable to simple text modifications,
Negative impact of randomness to generated output. , green-box-lv4, text width=31em] ] [Sentence-based
Sampling, green-box-lv3 [Mechanism: SemStamp [77], k-SemStamp [78].
Strengths: More resilient to small edits.
Weaknesses: Require more phases, resources and complex algorithms to
train; Limited scope: Works only when appropriately trained;
Susceptible to sentence-reordering attacks.
, green-box-lv4, text width=31em] ] ][WM During
LLM Training, orange-box-lv2 [Trigger-based
Watermarking, orange-box-lv3 [Mechanism: CodeMark [79], CoProtector [80], Hufu [81], WLM [82],
PLMmark [83].
Strengths: Controlled via trigger, Custom response.
Weaknesses: Ineffective without trigger, Potential misuse if discovered., orange-box-lv4, text width=31em] ] [Global
Watermarking, orange-box-lv3 [Mechanism: Distillation WM [84], RLWM [85], EmMark [86].
Strengths: Universal and integrated protection.
Weaknesses: Complicated to train, May affect model utility., orange-box-lv4, text width=31em] ] [LLM Architecture
Watermarking, orange-box-lv3 [Mechanism: Cross-Attention WM [87].
Strengths: Robust to synonym substitution and paraphrasing.
Weaknesses: Challenge in balancing WM robustness with the
quality of the generated text, Require fine-tuning, adding complexity to
the model’s deployment.
, orange-box-lv4, text width=31em] ] ] [Miscellaneous, violet-box-lv2 [Mixed Method
Watermarking, violet-box-lv3 [Mechanism: Duwak [88], Waterpool [89].
Strengths: Higher robustness, Improved detection.
Weaknesses: Increased computational overheads, Potential
interference between methods., violet-box-lv4, text width=31em] ] [Multiple-Output
Watermarking, violet-box-lv3 [Mechanism: WaterMax [90].
Strengths: Low quality impact.
Weaknesses: Slower generation, More compute., violet-box-lv4, text width=31em] ] [Prompt-Based
Watermarking, violet-box-lv3 [Mechanism: ModelShield [91].
Strengths: Adaptive, High text fidelity.
Weaknesses: Increased computational overheads, Vulnerable to
WM removal attack., violet-box-lv4, text width=31em] ] ] ]
@@ 3 Taxonomy of WMs and IP Checkers
To defend against model stealing attacks, recent works have highlighted the effectiveness and feasibility using WMs [9, 25, 50]. WMs embed a unique and detectable pattern into LLM outputs to trace and identify the origin of the text. A WM can be represented as a triple $(\mathcal{G},\mathcal{W},\mathcal{D})$, where $\mathcal{G}$ is the WM generator, which takes a prompt $x$ and a WM function $\mathcal{W}$ to produce a watermarked output $y^{wm}=\mathcal{G}(x,\mathcal{W})$, and $\mathcal{D}$ is a WM detection function, also known as an IP checker, to determine whether a given text is watermarked or not.
@@ 3.1 WM Generators
WM generators typically introduce perturbations to the probability distribution of LLMs during output generation. Depending on $\mathcal{W}$ and types of perturbations, WMs can be classified into several research directions, as shown in Fig. 3. First, WM during logits generation approaches partition tokens into green and red lists and add bias into logits to favor tokens from the green list. The tokens in the watermarked output are thus sampled with higher probability from the green list, which is determined by a pseudo-random generator seeded by the input prompt [9]. These methods are flexible and effective for tracking watermarked outputs but may impact the semantic meaning of the output and are vulnerable to removal attacks [26]. Second, WM during sampling process methods alter LLM text generation’s sampling process of next tokens or whole sentences without altering the underlying distribution of output. [25, 77]. While they are easy to integrate and allow for simple detection, they can be susceptible to attacks such as text modifications, randomness, and sentence reordering, especially in token-based methods. Third, WM during LLM training methods embed WMs by training LLMs with specific triggers, knowledge distillation, reinforcement learning, or by modifying the LLM architecture [85, 84, 80, 79, 87]. These methods offer inherent robustness but are often associated with implementation complexity and increased computational overheads. In addition, several miscellaneous WM approaches, such as mixed methods, multiple-output generation with selection, or prompt-based techniques, offer enhanced robustness and adaptability [90, 91, 89, 88]. However, these methods come with trade-offs, including higher computational costs and potential vulnerabilities to watermark removal attacks.

### v2

# SoK: Are Watermarks in LLMs Ready for Deployment?

@@ 1 Introduction
Large language models (LLMs), such as ChatGPT, Gemini, Claude, and Cohere [gemini, OpenAI, AnthropicClaude, Cohere], have demonstrated remarkable capabilities in text generation, machine translation, and knowledge understanding tasks [zhang2023prompting, zhang2023machine, xu2023paradigm, hu2023unbiased, kirchenbauer2023watermark]. They effectively mimic human writing behaviors and generate complex and coherent outputs from the input text, making it challenging to determine whether a text is authored by humans or generated by LLMs. Due to the high demands of computational resources and human efforts required for training LLMs [brown2020language, chen2021evaluating], these models are commonly offered as a service through application programming interfaces (APIs), typically requiring users to pay or subscribe [minthigpen, ibm]. Although users cannot access to the model weights or architectures of these commercial LLMs, this restriction does not ensure the safety of these models. Malicious actors can intentionally mimic cloud-hosted LLM behaviors to offer cheaper services [wallace2020imitation, xu2021student]. To conduct this service stealing, an adversary can query a set of inputs through an LLM’s API to retrieve the corresponding outputs. Then, the adversary uses these input-output data to effectively fine-tune their local model. When the number of queries is sufficient to gather enough input-output data within a particular domain, the adversary can steal the cloud-hosted LLM behaviors in that domain [googlebard, carlini2024stealing, aiattack]. Consequently, these concerns underscore significant risks regarding the intellectual property (IP) rights of the cloud-hosted proprietary LLMs [li2023protecting].
@@
To address such risks, service providers, i.e., the cloud, have employed many strategies including watermarks, encryption, limited model exposure via APIs, and differential privacy [kirchenbauer2023watermark, xue2022advparams, yu2021differentially]. Among them, watermarks (WMs) [kirchenbauer2023watermark, yoo2023advancing, liu2023unforgeable, christ2024undetectable, kuditipudi2023robust] have emerged as a practical tool for LLMs due to its ability to ensure traceability, protect IP, and detect misuse. Typically, WMs embed imperceptible patterns into the LLM outputs directly. These patterns can be used to trace the original text or determine whether the generated text from an LLM is watermarked, facilitating the detection of unauthorized use. To watermark outputs, the cloud can introduce bias into logits of token generations to favor a set of specific tokens or alter the token sampling process of LLMs. Once an WM is applied, the cloud can use IP checkers to determine if a set of outputs from a suspicious model is watermarked. This allows the cloud to assess whether the suspicious model has been trained or fine-tuned using the cloud’s watermarked outputs. Typically, IP checkers analyze the tokens and perform statistical tests to check if the values exceed certain thresholds, indicating that the cloud-hosted LLM has been imitated.
@@
WMs are effective in detecting IP violations. However, WMs can be susceptible to WM removal attacks [zhang2023watermarks, pang2024attacking] or spoofing attacks [pang2024attacking]. These attacks either require significant computational resources for additional training to paraphrase the generated text of cloud-hosted LLMs or demand numerous queries to learn the distribution of WMs. While several studies have surveyed WMs on LLMs [liu2024survey, liang2024watermarking, lalai2024intentions], no research has systematically and extensively explored the effectiveness of WMs on LLMs, particularly regarding the impact of WMs on LLM utility and IP checkers, under both attack and attack-free environments. In addition, the practicality of WMs has not been rigorously assessed. Therefore, in this work, we aim to develop a comprehensive systematization of WMs in LLMs, with the following contributions:
@@ 2.1 Large Language Models
Large language models (LLMs) [radford2019language, OpenAI, MicrosoftAzureAI, HuggingFaceMetaLLaMA] are advanced deep learning models, pre-trained on vast datasets with a large number of parameters, enabling them to perform a wide range of natural language processing (NLP) tasks, such as text classification, text and code generation, and question answering.
@@
In generation tasks, the model predicts each token in an autoregressive manner [radford2019language], similar to Eq. 1, to generate a complete text sequence.
The most widely used LLM architectures are based on transformers [vaswani2017attention] and incorporate a self-attention mechanism. This mechanism includes a stack of multi-head attention and feed-forward layers. Positional encoding with distinct matrices for different positions ensures the model retains sequence information without relying on recurrent networks. Based on transformer architectures, existing LLMs can be divided into three main categories (Fig. 1), as follows. (1) Encoder-only models, such as BERT [Devlin2019BERTPO], RoBERTa [liu2019roberta], XLNet [yang2019xlnet], or XLM [conneau2019cross], capture word context and are well-suited for tasks that require a deep understanding of input data, such as sentence classification, sentiment analysis, and information extraction. These models are trained bidirectionally, predicting masked words using both preceding and following words. (2) Decoder-only models, such as GPT and its family [Radford2018ImprovingLU], operate unidirectionally, predicting the next token based on context in an auto-regressive manner. They are designed for text generation tasks without a separate encoding phase. Lastly, (3) Encoder-decoder models such as T5 [raffel2020exploring], BART [DBLP:journals/corr/abs-1910-13461] or MASS [song2019mass] combine both encoder and decoder components for tasks that involve understanding and generating data, such as text summarization and translation.
@@ 2.2 Risks in LLMs
Due to the high demands of computational resources and human efforts required for training LLMs [brown2020language, chen2021evaluating], LLMs are commonly offered as a service through APIs, typically requiring users to pay or subscribe [minthigpen, ibm]. Table VI (Appendix) presents a comprehensive overview of popular LLM APIs, highlighting their features, advantages, and constraints. According to the Artificial Analysis leaderboard [leaderboard], OpenAI and Microsoft are currently the leading providers in the LLM API market.
Although users may not have access to the model weights or architectures via APIs, this restriction does not guarantee the model’s safety. In addition, as LLMs become more widespread, the risk of their misuse for malicious purposes is increasing. Potential misuses of LLMs include several significant threats, as follows. First, LLMs can be exploited to generate false information or carry out harmful actions, such as crafting phishing emails, hate speech, or biased content to manipulate public opinion or amplify harmful narratives [chen2023can]. Second, LLMs can generate deepfake data, making manipulated media more convincing and enabling sophisticated scams or damage to reputation [mitra2024world]. Third, adversaries can use LLMs to mimic individuals or organizations in text-based communications, facilitating theft, fraud, or financial crimes [chen2023combating]. Fourth, LLMs can assist in writing malicious code or crafting evasion techniques to bypass detection systems, posing a significant cybersecurity threat [motlagh2024large]. More importantly, malicious actors may intentionally exploit LLM behaviors by imitating cloud-hosted models to offer cheaper services or engage in model-stealing attacks, further amplifying the aforementioned risks. Such attacks not only threaten the security of proprietary models but also expose users to fraudulent or unethical services [wallace2020imitation, xu2021student].
Figure 2: Model Stealing Attacks [jovanovic2024watermark, dang2025delta].
@@ 2.3 Model Stealing Attacks
In model stealing attacks [jovanovic2024watermark, dang2025delta], adversaries aim to imitate the behavior of a cloud-hosted model denoted as $\theta$, by constructing their local model $\theta_{adv}$ through fine-tuning on input-output pairs obtained by querying the cloud-hosted model [zhang2023apmsa]. The goal is to bypass IP checkers while maintaining the adversary model’s utility, which is comparable to that of the original model. To launch the attacks (Fig. 2), adversaries start with a set of $N$ task-specific prompts $\{x_{i}\}_{i=1}^{N}$, sending them to the cloud to obtain outputs $\{y_{i}\}_{i=1}^{N}$. These pairs are used to fine-tune $\theta_{adv}$ to mimic the behaviors of $\theta$.
@@
[…]
Watermarking
for LLMs
WM During Logits Generation Mechanism: KGW [kirchenbauer2023watermark], SIR [liu2024semanticinvariantrobustwatermark], DiPmark [wu2024dipmark], SemaMark [ren-etal-2024-robust], Adaptive WM [liu2024adaptive], Unbiased WM [hu2024unbiased], TSW [pmlr-v235-huo24a], WatME [liang2024watme], GumbelSoft [fu2024gumbelsoft], MPAC [yoo2024advancing], UPV [liu2023unforgeable], Unigram [zhao2024provable], Topic-based WM [nemecek2024topic], REMARK LLM [zhang2024remark], SWEET [lee2023wrote], CTWL [wangtowards], EWD [lu2024entropy], X-SIR [he2024can], SW [fu2024watermarking], CodeIp [guan2024codeip], OW [wouters2023optimizing], Stylometric WM [niess2024stylometric], NS-WM [takezawa2023necessary]. Strengths: Flexibility and non-intrusive WM, Effectively track watermarked text. Weaknesses: Possible impact on semantic meaning, Vulnerable to removal attacks. WM During Sampling Process Token-based Sampling Mechanism: Undetectable WM [christ2024undetectable], EXP [kuditipudi2023robust], EDRW [golowich2024edit], EXPGumbel [aaronson2022watermarking], PDW [fairoze2023publicly], Bileve [zhou2024bileve], STA-1 [mao2024watermark], SynthID [dathathri2024scalable]. Strengths: Can be incorporated easily and less noticeably, Simple detection, checking the alignment between tokens and the sampling pattern. Weaknesses: Can be vulnerable to simple text modifications, Negative impact of randomness to generated output. Sentence-based Sampling Mechanism: SemStamp [hou2023semstamp], k-SemStamp [hou2024k]. Strengths: More resilient to small edits. Weaknesses: Require more phases, resources and complex algorithms to train; Limited scope: Works only when appropriately trained; Susceptible to sentence-reordering attacks. WM During LLM Training Trigger-based Watermarking Mechanism: CodeMark [sun2023codemark], CoProtector [sun2022coprotector], Hufu [xu2024hufu], WLM [guwatermarking], PLMmark [li2023plmmark]. Strengths: Controlled via trigger, Custom response. Weaknesses: Ineffective without trigger, Potential misuse if discovered. Global Watermarking Mechanism: Distillation WM [gu2023learnability], RLWM [xu2024learning], EmMark [zhang2024emmark]. Strengths: Universal and integrated protection. Weaknesses: Complicated to train, May affect model utility. LLM Architecture Watermarking Mechanism: Cross-Attention WM [baldassini2024cross]. Strengths: Robust to synonym substitution and paraphrasing. Weaknesses: Challenge in balancing WM robustness with the quality of the generated text, Require fine-tuning, adding complexity to the model’s deployment. Miscellaneous Mixed Method Watermarking Mechanism: Duwak [zhu2024duwak], Waterpool [huang2024waterpool]. Strengths: Higher robustness, Improved detection. Weaknesses: Increased computational overheads, Potential interference between methods. Multiple-Output Watermarking Mechanism: WaterMax [giboulot2024watermax]. Strengths: Low quality impact. Weaknesses: Slower generation, More compute. Prompt-Based Watermarking Mechanism: ModelShield [pang2024adaptive]. Strengths: Adaptive, High text fidelity. Weaknesses: Increased computational overheads, Vulnerable to WM removal attack.
@@ 3 Taxonomy of WMs and IP Checkers
To defend against model stealing attacks, recent works have highlighted the effectiveness and feasibility using WMs [kirchenbauer2023watermark, kuditipudi2023robust, liu2024semanticinvariantrobustwatermark]. WMs embed a unique and detectable pattern into LLM outputs to trace and identify the origin of the text. A WM can be represented as a triple $(\mathcal{G},\mathcal{W},\mathcal{D})$, where $\mathcal{G}$ is the WM generator, which takes a prompt $x$ and a WM function $\mathcal{W}$ to produce a watermarked output $y^{wm}=\mathcal{G}(x,\mathcal{W})$, and $\mathcal{D}$ is a WM detection function, also known as an IP checker, to determine whether a given text is watermarked or not.
@@ 3.1 WM Generators
WM generators typically introduce perturbations to the probability distribution of LLMs during output generation. Depending on $\mathcal{W}$ and types of perturbations, WMs can be classified into several research directions, as shown in Fig. 3. First, WM during logits generation approaches partition tokens into green and red lists and add bias into logits to favor tokens from the green list. The tokens in the watermarked output are thus sampled with higher probability from the green list, which is determined by a pseudo-random generator seeded by the input prompt [kirchenbauer2023watermark]. These methods are flexible and effective for tracking watermarked outputs but may impact the semantic meaning of the output and are vulnerable to removal attacks [zhang2023watermarks]. Second, WM during sampling process methods alter LLM text generation’s sampling process of next tokens or whole sentences without altering the underlying distribution of output. [kuditipudi2023robust, hou2023semstamp]. While they are easy to integrate and allow for simple detection, they can be susceptible to attacks such as text modifications, randomness, and sentence reordering, especially in token-based methods. Third, WM during LLM training methods embed WMs by training LLMs with specific triggers, knowledge distillation, reinforcement learning, or by modifying the LLM architecture [xu2024learning, gu2023learnability, sun2022coprotector, sun2023codemark, baldassini2024cross]. These methods offer inherent robustness but are often associated with implementation complexity and increased computational overheads. In addition, several miscellaneous WM approaches, such as mixed methods, multiple-output generation with selection, or prompt-based techniques, offer enhanced robustness and adaptability [giboulot2024watermax, pang2024adaptive, huang2024waterpool, zhu2024duwak]. However, these methods come with trade-offs, including higher computational costs and potential vulnerabilities to watermark removal attacks.